OpenStudy (anonymous):

find the range of f(x) =-x+4 for the domain {-3,-2,-1,1

OpenStudy (sidsiddhartha):

domain:x values of a function range:y values of a function just put those x values -3,-2,-1,1 one by one in the equation and u'll get y values which will be ur range

OpenStudy (sidsiddhartha):

like f(x)=-x+4 f(-3)=3+4=7 f(-2)=2+4=6 like this

OpenStudy (anonymous):

but the answer choices are -7-6-5-4 7,6,5,3 7,6,5,4 -7-6-5,3

OpenStudy (sidsiddhartha):

yes f(-3)=7 ,f(-2)=6,f(-1)=1+4=5,f(1)=-1+4=3 so {7,6,5,3}

OpenStudy (anonymous):

thanks so much

OpenStudy (sidsiddhartha):

:) u can click the best response :P